Shift summary for OWL shitf, June 9, 1999

Alexander Ledovskoy (ledovsko@ktev2)
Wed, 9 Jun 1999 07:10:16 -0500 (CDT)

Shift Summary for OWL shift, 9-June-99

On Shift: P. Toale
S. Ledovskoy

Also: Ed and Rick K.

Hours of beam delivered at all intensities: 7
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

2:00-5:30 Stream 3 debugging

========================================================================
Runs ended this shift:
Run 13455 - 1M events - BEAM_MUON with NM4AN at 400A
Run 13456 - 240 events - BEAM_832

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

3:20-5:30 Rick K and Sasha L traced Stream 3 problems.

========================================================================
Overview of shift:
E832 datataking interrupted by Stream 3 problems

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13455 muon run with NM4AN at 400A still going

00:30 Online split asks to mount QKMU03 on ktev23 tps3d2
Done. Tape drive starts blinking

00:45 Rick K. changed TOFF_L1SRC from 43.0 to 41.7 ns
in all trgmon*.nml This decision is based on
comparison of HC-RF time difference from online
histograms and from the scope.

13456 BEAM_832 started

13 1:05 CsI ch.1847 dropped its gain by 60%
Ed cycled the voltage for this channel
Up to this spill this channel has voltage off

1:10 Pipeline errors - Quick Reset
1:12 Pipeline errors - Reset with DBC
Intensity is 1.3E12 now
1:15 Pipeline errors

Divider for 1847 needed to be swapped (later).
We are running with bad 1847

1:25 Repeater in laser crate is causing problems - Cycled.

Stream 2 event size (l3#summary) has abnormal peak
at high values (see logbook)

Rick K. Reduced accidental trigger by factor of 10
It was 10 time larger then expected based from
the muon running

1:48 ktev1 and ktev2 are not processing any events now.
However, triggers are going to all planes
stop this run

Scaler:LIVE_TIME=0 alarm goes off when we are not
running. (e-mailed to Theo)

1:50 Fresh start

2:05 No beam. Ch13 says "Beam to users"

13457 BEAM_832 started

2:13 Beam is back.
2:20 No beam again

2:35 Control room called: they will tune down intensity
to about 1E12 to work on beam

ktev4 is the only live node (again).
Fresh start again
we reset all DDDs (buttons on the back)

13458 BEAM_832 started
We aborted previous run instead of stopping it. That
could be a problem.

2:45 Online split tape drive (with QKMU03) stop blinking

Fresh start again. Starting new run without
writing on tapes.

On a positive note, the intensity is still at 1.6E12

13459 BEAM_832 started (no tapes)
Nothing goes to memory
Rick K. declares Stream 3 dead.

3:20 After checking things upstairs (did not find anything
wroing there) we requested an access.

It took a while to find that repeators for stream 3
downstairs had to be cycled to fix DAQ problems

5:30 Access is over. Beam is back (0.75E12).

13463 BEAM_832 started. We OVERWRITING tapes because
we aborted runs with fresh start several times.

CA channel 3 has no hits

6:20 wanted to do a gas checklist but could not find
a binder with it. Used an old gas checklist as
a guideline: everything is OK

49 6:40 HCC buffer is full (second time during this shift)

6:50 Pipeline errors happend twice during last 10 min


========================================================================
Follow up Notes on Previous Shift

- Fixed ktevana equivalence bug for VV' and RC trigger sources.
On-line monitor plots OK.

- CA and MA ADC now extend up to about 200 counts instead of 150 counts
due to reducing the DC offset into the ADC.

========================================================================
Items to Follow up on Next Shift

1. Accidental 90 for hadron runs (R.Colman is contacted)
2. Scaler:Live_TIME alarm (Theo is contacted)
3. CsI-1847 divider swap
4. CA channel has no hits (see trg#vetohit, camon#summary)